OpenStudy (anonymous):

Solve the equation 2/3k + 6 = −12 for k. Please show your work and solution in the space provided. PLEASE HELP ME!!!!!!

OpenStudy (anonymous):

Have you gotten any help with this yet?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

So let's get started then :)

OpenStudy (anonymous):

First, we need to isolate your variable (k) on one side. Doesn't matter what side you choose, but I have a habbit of putting my variables on the left.

OpenStudy (anonymous):

I do too

OpenStudy (anonymous):

If we do this, then you get 2/3k=-12-6 or 2/3k=-18

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

In order to get the 2/3 to drop out on the side with the variable, we need to multiply 2/3 by its reciprocal which is 3/2. We do this to both sides.

OpenStudy (anonymous):

Set up like this: (3/2) (2/3k)= -18 (3/2) Are you with me so far?

OpenStudy (anonymous):

Kinda... I've never worked out a problem like that

OpenStudy (anonymous):

We just have to remember that what we do to one side, we do other :) So, when you multiply the fraction on the left, you get 6/6 * k which is 1*k. 1*k is just k.

OpenStudy (anonymous):

Now to the right side. You h ave -18 * (3/2). It is really (-18 * 3)/2.

OpenStudy (anonymous):

So, it is -54/2 which gives you -27.

OpenStudy (anonymous):

Okay I think i get it. Thanks for the help

OpenStudy (anonymous):

You can always double check by substituting your value for k and solcing the equation. Glad I could help :)

OpenStudy (anonymous):

Sorry- solving for k.